In current electricity
power (P) = I^2R and also
power (P) = V^2/R
Is power proportional to R or inversly proportional to R ??

Grade:12

2 Answers

akhil mahajan
13 years ago

P= I2R when V is constant

and P=V2/R when I is constant..

Vikas TU
13 years ago

P = I22R .......(1)

p = V22/R .........(2)

from   (1)  and (2)

I22R = V22/R 

(IR)2 = (V)2

V = IR 

R = V/I         :. [P = VI]

                        But here VI is in ratio form

so, P will be inversely propirtional to R

:-)